[Brahms-l] Shift report 20040302 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Tue Mar 02 2004 - 16:01:50 EST
------------------------------------------------------------
Date: Tue Mar  2 16:01:50 2004, Shift: 08:00-16:00
Supervisor: HB, with: PJ, FH
Log book#15 pages 72 - 72

Shift Summary:
------------------------------------------------------------
1. Beam lost unexpectedly at 11:15. New store again 14:22, but with rather low luminosity.

2. All runs: FS 8deg. full B  MRS 40deg. 1/2B

10726,10727,10728,10729,10730. Adding to a total tr2 232k and tr3 650k

3. Fine. except that the DAQ Acted strange during run 10730 (it's now marked as bad). So we restarted the IoServer and RunServer.

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Mar 2 16:02:55 2004

This archive was generated by hypermail 2.1.8 : Tue Mar 02 2004 - 16:03:18 EST